Charlie Company

The Balance Sheet and Income Statement of Charlie Company as of and for the twelve months ended December 31, 2018 shows the following information (in thousands):

Opening (Prior Period) and Ending balances

Accounts Receivable Opening $4,155, Ending. $5,155

Accounts Payable Opening $5,200, Ending $5,750

Inventory Opening $2,150, Ending 4,150

Sales on Credit (12months) $42,450

Cost of Goods Sold (COGS) (12 months) 31,500

QUESTION: Using this information, how many days were in the Inventory Conversion Period for the year (ROUND UP TO WHOLE DAY IF NECESSARY)?

A 36 days

B 33 days

C 40 days

D None of the above
Answer #1

Inventory turnover ratio = COGS/ Average inventory

Average inventory = 2150+ 4150 / 2 = 3150

= 31500/3150 = 10 times

Inventory conversion period = 360 days / 10 = 36 days

Option A.